Prova scritta di AM4 del 6/9/2004 - Appello X Soluzioni

1) 4 + \frac{2}{\sqrt{3}} log(2 + \sqrt{3})

2) u(x, t) = \frac{1}{2π} ∫_{{∞ \atop −∞}} \frac{sen ξ}{ξ} e^{−ξ^{2}t} cos ξ(x − t) dξ

3) \frac{π^{4}}{90}